OpenStudy (anonymous):

Write an equation of the line in point-slope form, that passes through the two given points Points (-2,15), (9,-18) Answers are Y-2=-3(x-15) Y-2=-1/3(x+15) Y-15=-3(x+2) Y-15=-1/3(x+2) Thanks

OpenStudy (vshiroky):

y2-y1/x2-x1 will give you your slope. then do y-y1=slope(x-x1)

OpenStudy (a1234):

Slope formula- (y2 - y1)/(x2 - x1) (-18 - 15)/(9 + 2) -33/11 = -3 Point-slope form: y - y1 = m(x - x1) y - 15 = -3(x + 2) y - 15 = -3x - 6 y = -3x + 9

OpenStudy (a1234):

C.